OpenStudy (anonymous):

What is the length of side x? If necessary, round your answer to two decimal places. the hypotenuse is 26 units and the long leg is 18.

OpenStudy (anonymous):

18.76

OpenStudy (anonymous):

as we now : hypotenuse^2 = perpendicular^2 + base^2 (26unit)^2 = (18unit)^2 + (base)^2 (26unit)^2 - (18unit)^2 = x^2 as we know a^2-b^2 is (a+b)(a-b) so we will get (26+18)(26-18) = x^2 352 = x^2 root of 352 = x that is 18,76 PLS GIVE ME MEDAL IF U LIKE IT

OpenStudy (anonymous):

THANKS FOR GIVING ME MEDAL

OpenStudy (anonymous):

your welcome